### v1

# Superdiffusion and antidiffusion in an aligned active suspension

Abstract
We show theoretically that an imposed uniaxial anisotropy leads to new universality classes for the dynamics of active particles suspended in a viscous fluid. In the homogeneous state, their concentration relaxes superdiffusively, stirred by the long-ranged flows generated by its own fluctuations, as confirmed by our numerical simulations. Increasing activity leads to an anisotropic diffusive instability, driven by an active contribution to the particle current proportional to the local curvature of the suspension velocity profile.
**footnotetext: These authors contributed equally to this work.
Active Matter, that is, living materials and their imitations, is made up of components that continually extract mechanical work from a fuel supply [1, 2]. The dynamical equations governing active materials differ from their passive counterparts through physical fluxes arising from maintained chemical forces [3]. In the cases of greatest interest, these fluxes dominate at large length scales, resulting in dynamical and statistical properties qualitatively different from those of a time-reversal-invariant system with the same spatial symmetries. Much insight has emerged recently [4, 5, 6, 7] from a focus on active currents and stresses built from scalar fields [8] rather than alignment or flocking. In this Letter we study such scalar active matter in a permanently anisotropic momentum-conserving fluid, described, like active [9, 6] or passive [10] model H, by concentration and hydrodynamic velocity fields, but with a preferred axis. We discuss possible experimental realizations, e.g., Fig. 1, towards the end of the paper.
Figure 1: (a) Schematic realization of (b) tumbling but aligned swimmers, in the form of bacteria homogeneously dispersed in a stiff nematic liquid crystal that aligns their force dipoles along its fixed $\hat{\bf z}$ axis. (c) Far-field fluid velocity field due to a swimmer.
Here are our main results. The self-advection of concentration fluctuations by the long-ranged active flows they generate leads to a dynamic exponent $z=d/2$ for $d\leq 4$: inhomogeneities on a length scale $L$ relax on a timescale $\propto L^{3/2}$ in $d=3$. Our numerical studies of Brownian force dipoles in a Stokesian fluid confirm this asymptotic superdiffusive scaling, and uncover an early-time ballistic dynamics that also emerges from our theory. In addition, we point out the existence of an active contribution to solute currents proportional to the curvature in the profile of the hydrodynamic velocity field (Fig.˜2), that can drive an anisotropic diffusive instability, with a critical point at which the hydrodynamic interaction is strongly relevant. We thus uncover two hitherto unknown universality classes, describing the homogeneous phase and the onset of phase separation of uniaxial active model H.
We now show how these results were obtained. We work with an active suspension, with solute concentration field $c({\bf r},t)$ and joint velocity field ${\bf u}({\bf r},t)$, as functions of position $\bf r$} and time $t$, and a macroscopic anisotropy defined by a distinguished $\hat{\bf z}$ direction, with $\hat{\bf z}\to-\hat{\bf z}$ symmetry and isotropy in the $\perp$ plane transverse to $\hat{\bf z}$. We enforce overall incompressibility, i.e., total density $\rho=$ constant, as appropriate for the slow flows of interest here, so that $\nabla\cdot{\bf u}=0$. The imposed, rather than spontaneous, alignment means that $c$ and $\bf u$} are the only slow variables, obeying the conservation laws $\partial_{t}c=-\nabla\cdot{\bf J}$ for number and $\partial_{t}(\rho{\bf u})=\nabla\cdot{\bm{\sigma}}$ for momentum, with current $\bf J$} and stress tensor $\bm{\sigma}$. To build our equations of motion, we need only ask what contributions to $\bf J$} and $\bm{\sigma}$, constructed from $c$ and $\bf u$} and their gradients, can arise in the presence of anisotropy and activity, that compete with those already accounted for in (isotropic) active model H [9, 6]. A little reflection will show that there are precisely two contributions, at linear order in the fields:
$\bm{\sigma}^{a}=-W\hat{\bf z}\hat{\bf z}c({\bf r},t)$ (1)
to the stress, linear in the concentration, and
$\mathbf{J}_{\perp}^{\mathbf{u}}=$ $\;a_{1}\nabla_{\perp}\partial_{z}u_{z}+(a_{2}\nabla_{\perp}^{2}+a_{3}\partial_{z}^{2}+a_{4}\nabla_{\perp}\nabla_{\perp}\cdot)\mathbf{u}_{\perp},$
$J_{z}^{\mathbf{u}}=$ $\;b_{1}\partial_{z}^{2}u_{z}+b_{2}\partial_{z}\nabla_{\perp}\cdot\mathbf{u}_{\perp}+b_{3}\nabla_{\perp}^{2}u_{z},$ (2)
which we term Flow-Induced Motility or FIM, to the current. Galilean invariance, incompressibility, and symmetry under inversion of and rotation about $\hat{\bf z}$ rule out any contribution (apart from $c{\bf u}$) of $\bf u$} to the particle current at lower gradient order than (Superdiffusion and antidiffusion in an aligned active suspension). Eqs. (1) and (Superdiffusion and antidiffusion in an aligned active suspension) lead to the equations of motion
$(\partial_{t}+{\bf u}\cdot\nabla)c=-(a\nabla_{\perp}^{2}+b\partial_{z}^{2})\partial_{z}u_{z}$
$+(D_{\perp}\nabla_{\perp}^{2}+D_{z}\partial_{z}^{2})c+\nabla\cdot\bf f+$... (3)
and
$\rho(\partial_{t}+{\bf u}\cdot\nabla){\bf u}=\eta\nabla^{2}{\bf u}-{\nabla\mathcal{P}}-W\hat{\bf z}\partial_{z}c+$... (4)
where $a,b$ are combinations of the $\{a_{i}\},\{b_{i}\}$ in Superdiffusion and antidiffusion in an aligned active suspension, the ellipsis denotes terms arising in [9, 6] or trivially anisotropic variants, $D_{\perp}$ and $D_{z}$ are bare diffusivities and $\eta$ is the shear viscosity, taken for simplicity to be isotropic, and the pressure $\mathcal{P}$ enforces incompressibility. In Superdiffusion and antidiffusion in an aligned active suspension, we have allowed for number-conserving fluctuations in the form of a Gaussian random current $\mathbf{f}(\mathbf{r},t)$, white in space and time, with strengths $_{z}, NN_{\perp}$ respectively along and transverse to $\hat{\bf z}$. In Eq.˜4 $W$ is positive (negative) for extensile or pusher (contractile or puller) swimmers. The presence of active forcing linear in $\nabla c$, impossible in the isotropic theory [6], is the central feature of Eq.˜4. Superdiffusion and antidiffusion in an aligned active suspension and 4 constitute Uniaxial Active Model H. For the remainder of this paper, we will disregard inertia by setting the left-hand side of Eq.˜4 to zero, a valid approximation for describing microbial suspensions. We will then see that the velocity field $\bf u$} qualitatively changes the concentration dynamics through (Superdiffusion and antidiffusion in an aligned active suspension) and (1).
We begin by neglecting advection ${\bf u}\cdot\nabla c$ in Superdiffusion and antidiffusion in an aligned active suspension and studying the consequences of FIM for the effective linearized dynamics of $c$. Defining spatial Fourier components $c_{\bf k}(t)=\int_{\bf r}c({\bf r},t)e^{-i{\bf k}\cdot{\bf r}}$, and similarly for $\bf u$}, Eq.˜4 and incompressibility imply ${\bf u}_{\bf k}={\bf B}_{\bf k}c_{\bf k}$, where
${\bf B}_{\bf k}=-iW\frac{\bm{\Pi}_{\bf k}}{\eta k^{2}}\cdot\hat{\bf z}k_{z},\,\,\mbox{and}\,\,\bm{\Pi}_{\bf k}=\textsf{{I}}-\hat{\bf k}\hat{\bf k}$ (5)
projects transverse to $\bf k$}. Inserting the result in the Fourier-transformed Superdiffusion and antidiffusion in an aligned active suspension and neglecting advection gives $\partial_{t}{c_{k}}=-D(\theta)k^{2}c_{k}$ where $\theta$ is the angle between the wavevector $\bf k$} and the $\hat{\bf z}$ axis, with
$D(\theta)=\left(D_{z}-\frac{bW}{4\eta}\sin^{2}2\theta\right)\cos^{2}\theta$
$+\left(D_{\perp}-\frac{aW}{4\eta}\sin^{2}2\theta\right)\sin^{2}\theta$ . (6)
Some remarks are in order at this stage. First, our decision to invoke FIM stands vindicated as its effect enters Superdiffusion and antidiffusion in an aligned active suspension at the same order in wavenumber as the bare diffusive dynamics. Further, the diffusivity shifts $aW/\eta$, $bW/\eta$ are quadratic in activity, relying on active contributions to both Superdiffusion and antidiffusion in an aligned active suspension and 4, in contrast to the familiar scenario, see, e.g., [11, 12, 13], of interplay between a passive and an active term.
Figure 2: Left circles represent an initial isotropic configuration of concentration. Red arrows indicate the directions of $\bf u$} (a) when $\partial_{\perp}\partial_{z}u_{z}>0$, (b) when $\partial_{z}\partial_{\perp}u_{\perp}>0$, where $\perp=x$ or $y$, and the right blue blobs are the concentration distribution after advection by the red-arrow flows.
Symmetry allows either sign for $a,b$, but Stokesian-swimmer physics shows $aW$ and $bW$ are always positive: Fig.˜2 shows flows with (a) $\partial_{\perp}\partial_{z}u_{z}>0$ and (b) $\partial_{z}\partial_{\perp}u_{\perp}>0$. If $a,b>0$, then Superdiffusion and antidiffusion in an aligned active suspension tells us these give rise respectively to currents $J^{u}_{\perp}>0$ (rightward) and $J^{u}_{z}<0$ (downward). Thus, the blobs in (a) and (b) move with their centres of drag ahead and astern respectively. This means (see, e.g., [14]) that with respect to the $\hat{\bf z}$ direction both are pushers, i.e., $W>0$. Similarly, $W<0$ for $a,b<0$. Therefore, aW,bW>0aW,\ bW>0, so that FIM, via Superdiffusion and antidiffusion in an aligned active suspension, always reduces the diffusivity. For large enough $a$ or $b$ the diffusivity will first hit zero for a cone of directions determined by Superdiffusion and antidiffusion in an aligned active suspension. Beyond this threshold $D(\theta)<0$ for a range of $\theta$ – the anti-diffusion alluded to in the title – i.e., a linear instability to small-wavenumber concentration fluctuations. Active hydrodynamics thus leads to phase separation without attractive interactions, with structural anisotropy playing the crucial role of orienting the persistent large-scale flows of the swimmers. The condensation can take place in an infinite system, with no need to invoke boundaries [15]. This discussion of course presumes an appreciable magnitude for $a$ and $b$. In the End Matter, we estimated them to be of order $(W\tau/\eta R)c_{0}$ where $\tau$ is the structural relaxation time of the suspension and $R$ the particle size. See End Matter for a discussion on possible effects of the hydrodynamic interaction at the putative critical point of this flow-induced condensation.
Note that even when the system is confined by walls to a 2D fluid film [16], the active currents generated by the averaged flow field along the confinement direction still contribute diffusion-like terms to the continuity equation Superdiffusion and antidiffusion in an aligned active suspension. This results in the possibility of condensation through the same mechanism [17]. Furthermore, although we assumed instantaneous steady Stokesian dynamics above, introducing viscous retardation for the velocity field – i.e., considering the unsteady Stokes equation – still leads to condensation with precisely the same diffusivity shifts as obtained earlier [17].
We focus now on the homogeneous suspension, for which it suffices to study (Superdiffusion and antidiffusion in an aligned active suspension) and (4) with $a,b=0$, and examine the effects of advection. We will proceed as usual [10, 18] by first establishing the scaling properties of the linear theory, and then testing the stability of the resulting Gaussian fixed point with respect to inclusion of the advective nonlinearity ${\bf u}\cdot\nabla c$. Rescaling ${\bf r}\to b{\bf r}$,   $t\rightarrow b^{z}t$,   $c\to b^{\chi}c$,   ${\bf u}\to b^{\Xi}{\bf u},\,\mathcal{P}\to b^{\psi}\mathcal{P}$, we find the Gaussian fixed point $z=2$, $\chi=-d/2$ is stable against the advective nonlinearity for $d>4$. For $d<4$ the large-scale long-time properties of the homogeneous phase of an anisotropic active suspension are governed by a nontrivial stable fixed point. We do not implement a dynamic renormalization-group to obtain the scaling properties for $d<4$. As Galilean invariance ensures the absence of fluctuation corrections to the vertex of the advective nonlinearity, it suffices to carry out a one-loop self-consistent calculation for the correlation function and propagator [19, 20].
We define the propagator $G({\bf r},t)\equiv\frac{\delta\langle c({\bf r},t)\rangle}{\delta h({\bf 0},0)}\bigg\rvert_{h=0}$ for $t>0$ and 0 for $t<0$, where $h({\bf r},t)$ is a source field added to the right-hand side of Superdiffusion and antidiffusion in an aligned active suspension, and the correlator $C({\bf r},t)=\langle c({\bf 0},0)c({\bf r},t)\rangle$, where the angle brackets are an average over the noise $\bf f$} in Superdiffusion and antidiffusion in an aligned active suspension. We work with the space-time Fourier transforms $G_{{\bf k}\omega}\equiv\int_{{\bf r},t}G({\bf r},t)e^{-i{\bf k}\cdot{\bf r}+i\omega t}\equiv\int_{t}G_{{\bf k}}(t)e^{i\omega t}$, and similarly $C_{{\bf k}\omega}$ and $C_{{\bf k}}(t)$. The one-loop self-consistent method for evaluating $G$ and $C$ [19, 20] consists in solving Superdiffusion and antidiffusion in an aligned active suspension iteratively by standard methods (see, e.g., [18, 21]), giving
$G^{-1}_{{\bf k}\omega}=G^{-1}_{0{\bf k}\omega}-\Sigma_{{\bf k}\omega}$ (7)
where $G_{0{\bf k}\omega}=(-i\omega+D_{\perp}k_{\perp}^{2}+D_{z}k_{z}^{2})^{-1}$ is the bare propagator,
$\Sigma_{{\bf k}\omega}=-\lambda^{2}\int\frac{d^{d}q}{(2\pi)^{d}}dt|V_{\bf kq}|^{2}C_{{\bf k}-{\bf q}}(t)G_{{\bf q}}(t)e^{i\omega t}$ (8)
the self-energy, and
$V_{\bf kq}=\frac{W}{2\eta}{\bf k}\cdot\left(\bm{\Pi}_{\bf q}\frac{q_{z}}{q^{2}}+\bm{\Pi}_{{\bf k}-{\bf q}}\frac{k_{z}-q_{z}}{|{\bf k}-{\bf q}|^{2}}\right)\cdot\hat{\bf z}$ (9)
the symmetrized vertex. The correlator is expressed in terms of $G$ and the renormalized noise strength $N_{{\bf k}\omega}$ as
$C_{{\bf k}\omega}=|G_{{\bf k}\omega}|^{2}N_{{\bf k}\omega},$ (10)
where
$N_{{\bf k}\omega}=N_{\perp}k_{\perp}^{2}+N_{z}k_{z}^{2}+\int\frac{d^{d}q}{(2\pi)^{d}}dte^{i\omega t}|V_{\bf kq}|^{2}C_{{\bf k}-{\bf q}}(t)C_{{\bf q}}(t).$ (11)
Self-consistency is enforced by the use of the full, not bare, propagator and correlator in the integrals in Eqs.˜8 and 11. To extract the scaling exponents $\chi,z$ it suffices to make the dynamical scaling ansatz $G_{{\bf k}}(t)=f(\Gamma k^{z}t),\,C_{{\bf k}}(t)=Ak^{-(d+2\chi)}g({\Gamma k^{z}}t)$, with coefficients $A$ and $\Gamma$ for the static correlator and the relaxation rate. We approximate the scaling functions $f(x)$ and $g(x)$ by $e^{-x}$, i.e., a single relaxation time for each wavenumber or equivalently a Lorentzian line-shape in the frequency domain, which is acceptable if the aim is only to identify scaling exponents. For $d<4$, relaxation at small $k$ is dominated by the self-energy, so we identify $\Sigma_{{\bf k}\omega}=\Gamma k^{z}$ and $N_{{\bf k}\omega}$ by the integral in Eq.˜11. Power-counting on Eq.˜8 can readily be seen to yield one condition on the two exponents, namely,
$\chi+z=0.$ (12)
Eq.˜12, which we obtained here by assuming the dominance of the fluctuation correction in the self-consistent calculation, is precisely the result that would emerge if we invoked Galilean invariance to keep the coefficient of ${\bf u}\cdot\nabla c$ in Superdiffusion and antidiffusion in an aligned active suspension fixed under rescaling. Scaling all wavevectors by $k$ itself and inserting the scaling ansatz allows us to rewrite Eqs.˜8 and 11 as
$\frac{2\Gamma^{2}}{A}$ $=\int\frac{d^{d}q}{(2\pi)^{d}}|V_{\hat{\bf k}\hat{\bf q}}|^{2}\frac{1}{q^{2}}\frac{1}{|\hat{\bf k}-{\bf q}|^{d+2\chi}}\frac{1}{q^{z}}$
$\quad\asymp\tilde{K}_{d}\frac{1-\Lambda^{-(2+2\chi+z)}}{2+2\chi+z},$ (13)
and
$\frac{2\Gamma^{2}}{A}$ $=\int\frac{d^{d}q}{(2\pi)^{d}}|V_{\hat{\bf k}\hat{\bf q}}|^{2}\frac{1}{q^{2}}\frac{1}{|\hat{\bf k}-{\bf q}|^{d+2\chi}}\frac{1}{q^{d+2\chi}}\frac{2}{|\hat{\bf k}-{\bf q}|^{z}+q^{z}}$
$\quad\asymp\tilde{K}_{d}\frac{1-\Lambda^{-(2+3\chi+d)}}{2+3\chi+d},$ (14)
where we retain the symbol $\bf q$} for the rescaled internal wavevector and have scaled the powers of $q$ out of the vertex. In Superdiffusion and antidiffusion in an aligned active suspension and Superdiffusion and antidiffusion in an aligned active suspension, $\Lambda$ is the ultraviolet cutoff, which we can take to $\infty$ for $d<4$, $\tilde{K}_{d}$ is an angular integral, and the asymptotic equality ($\asymp$) in each case is the large-$q$ contribution which dominates for $d\to 4$ (and turns into the logarithmic ultraviolet divergence for $d=4$) [19, 20]. Equating these for $d=4-\epsilon$, we see that
$d+2\chi=0,$ (15)
meaning that within our isotropic scaling ansatz the spatial rescaling properties of $c$ are always those of the Gaussian theory. The dynamical scaling properties are however nontrivial: combining Eq.˜15 with Eq.˜12 implies a superdiffusive dynamic exponent
$z=2-\epsilon/2,$ (16)
i.e., time $\sim$ length3/2 in $d=3$, or mean-square displacement $\sim t^{4/3}$.
Two arguments, one intuitive, the other formal, help us understand the self-consistent scaling results and their limitations. Consider a suspension of force dipoles of identical strength $W$, distributed randomly and uniformly with mean number density $c_{0}$, uniaxially aligned, in a fluid with viscosity $\eta$. The mean flow is zero as the stress is uniform. A local concentration fluctuation gives rise to a flow field of magnitude $\sim r^{-(d-1)}W/\eta$ a distance $r$ away. Assuming uncorrelated fluctuations and adding the squared flow fields incoherently gives a velocity variance $U^{2}_{RMS}\sim c_{0}(W/\eta)^{2}L^{-d+2}$ on a length scale $L$. The corresponding timescale is $\tau=L/U_{RMS}\sim(c_{0}^{1/2}\eta/W)L^{d/2}$, which trumps diffusion if $d<4$, reproducing both the dynamic exponent and the upper critical dimension that we reported above. Of course this argument can’t tell us that it was correct to treat the swimmers as independent. Our self-consistent calculation yielded both $z=d/2$ and $\chi=-d/2$, the latter implying the absence of positional correlations at large length scales, unlike in sedimentation, where the hydrodynamic interaction drastically alters [22, 23, 24] the independent-particle calculation [25].
More formally, if we set $a=b=0$ in Superdiffusion and antidiffusion in an aligned active suspension, and choose the bare noise and diffusion to obey $N_{z}/N_{\perp}=D_{z}/D_{\perp}$, then the steady-state solution of the Fokker-Planck equation for the probability distribution functional for the field $c$ in the linear theory (i.e., ignoring advection) is $P_{s}[c]\propto\exp[-\mbox{const.}\int_{\bf x}(\delta c)^{2}]$ so that the static correlator $C_{\bf k}(0)$ is a constant, i.e., independent of $\bf k$}. Does this result survive the inclusion of advection? It does, because the functional divergence $\int_{\bf x}{\delta}J/{\delta c({\bf x})}$ of the probability current $J={\bf u}\cdot\nabla cP_{s}[c]$ that the advective term ${\bf u}\cdot\nabla c$ induces, given the distribution $P_{s}$, is identically zero (see End Matter). Thus $P_{s}[c]\propto\exp[-\mbox{const}\int_{\bf x}(\delta c)^{2}]$ is an exact steady-state solution to the Fokker-Planck equation even when the advective nonlinearity is included. Then $\chi=-d/2$ and therefore, for $d\leq 4$, as Galilean invariance implies $\chi+z=0$, the dynamic exponent $z=d/2$. If $N_{z}/N_{\perp}\neq D_{z}/D_{\perp}$, these arguments do not apply, and a more complicated scenario emerges which we discuss briefly in End Matter.
Figure 3: Mean-Square-Displacement(MSD) of the particles, after substracting diffusive contributions, in log-log scale, (inset) in linear-linear scale after dividing by corresponding superdiffusive scaling, for different Péclet numbers.
We now test our field-theoretic predictions against a numerical simulation of Brownian point force dipoles of strength $W$, situated at positions $\{\mathbf{r}_{\alpha}(t),\,\alpha=1,...N\}$, aligned in the $\hat{\bf z}$ direction, suspended in a Stokesian fluid with viscosity $\eta$ and velocity field $\mathbf{u}(\mathbf{r})$, and obey the advected Brownian equation of motion
$\dot{\mathbf{r}}_{\alpha}=\mathbf{u}(\mathbf{r}_{\alpha})+\sqrt{2D}\;\bm{\zeta}_{\alpha}(t),$ (17)
where $D$ is the diffusivity and $\bm{\zeta}(t)$ is isotropic unit Gaussian white noise. We work with a 3D domain with periodic boundary conditions. Scaling lengths by the simulation grid size $\ell$ and times by $\ell^{2}/D$, the non-dimensionalized Eulerian velocity field due to the force dipoles is $\mathbf{u}(\mathbf{r})=\mathrm{Pe}\sum_{\mathbf{k}}\sum_{\alpha}{ie^{i\mathbf{k}\cdot(\mathbf{r-r_{\alpha}})}}\big(k_{z}^{2}\mathbf{k_{\perp}}-k_{\perp}^{2}k_{z}\hat{\bf{z}}\big)/{k^{4}}$, with Péclet number $\mathrm{Pe}=W/\eta D\ell$. The velocity field $\mathbf{u}$ evaluated on the simulation grid is interpolated to the $\{\mathbf{r}_{\alpha}(t)\}$ using a 4-point immersed boundary kernel [26].
We consider a cubic box of size $200^{3}$, seed 4000 particles uniformly throughout the system and run the simulation for $10^{6}$ iterations with a time step of $\delta t=5\times 10^{-4}$, using the Euler-Maruyama algorithm (see Supplemental Material [17] for further details).
Fig.˜3 displays the excess Mean-Square Displacement (MSD) with respect to ordinary diffusion (hereafter the “filtered MSD”), $\Delta\langle[\Delta\mathbf{r}(t)]^{2}\rangle\equiv\langle|\mathbf{r}(t)-\mathbf{r}(0)|^{2}\rangle-6t$ in the statistical steady state. We see an important feature apart from the asymptotic long-time superdiffusion, namely, an early-time ballistic regime, which we now show is also predicted by our theory. The filtered MSD of the particles is determined entirely by the velocity field generated by the force dipoles: $\Delta\langle[\Delta\mathbf{r}(t)]^{2}\rangle=\int_{s=0}^{t}ds\int_{s^{\prime}=0}^{t}ds^{\prime}\;\big\langle\mathbf{u}(\mathbf{r}_{\alpha}(s),s)\cdot\mathbf{u}(\mathbf{r}_{\alpha}(s^{\prime}),s^{\prime})$. Eq. (4) tells us that the velocity correlator in the integrand has a well-defined $s\to s^{\prime}$ limit which, for $t\to 0$, constitutes the coefficient of $t^{2}$, the leading short-time behavior of the filtered MSD [17]. Fig.˜3 (inset) shows the filtered MSD divided by the proposed asymptotic superdiffusive contribution to the MSD (i.e., $L^{2}\sim\tau^{\frac{4}{3}}$ for 3D), revealing a plateau that persists over a long time, confirming the prediction (16). Note that the noise-averaged particle-phase velocity field due to Eq.˜17 is divergenceless and therefore cannot generate the physics of FIM (Superdiffusion and antidiffusion in an aligned active suspension) and the diffusive instability (Superdiffusion and antidiffusion in an aligned active suspension). The latter would require motion across streamlines of $\bf u$}, as would arise in the presence of interparticle interactions via a pair potential.
To summarise: we have shown that the large-scale, long-time properties of a suspension of active particles in a viscous fluid with a built-in preferred axis lie in universality classes distinct from those of an isotropic active suspension with the same slow variables. Unlike in [6], activity enters even in the linearized dynamics of the concentration and hydrodynamic velocity. In the homogeneous phase, we predict superdiffusive relaxation of the concentration, with dynamic exponent $z=d/2$ in dimension $d<4$, i.e., time $\propto$ (length)3/2 in 3D. Our 3D numerical experiments with aligned Brownian force dipoles in a viscous fluid show a crossover from an early-time ballistic to the asymptotic superdiffusive behavior, confirming our proposed scaling. We have identified an active-particle current induced by curvature in the suspension velocity field, not discussed earlier in the literature, that drives phase separation through hydrodynamic flow. The underlying mechanism is fundamentally multi-particle in nature and thus inaccessible to the treatment of [14]. By contrast with the flow-induced phase separation of [15], it does not rely on sample boundaries. Possible physical realizations on which to test our theory include microbes dispersed homogeneously in a stiff nematic [27], or subjected to polarized [28] or directional [29, 30, 31] light fields, or colloidal suspensions in a uniform electric field [32, 33, 34], in which concentration fluctuations are force dipoles.
Acknowledgments— We thank J K Bhattacharjee, R Golestanian, A Maitra and P Popli for valuable discussions. SKS acknowledges computational support from Supercomputer Education and Research Centre, IISc. SR acknowledges support from a JC Bose Fellowship of the ANRF, India, and thanks the Isaac Newton Institute for Mathematical Sciences for support and hospitality during the programmes “New statistical physics in living matter: non equilibrium states under adaptive control” and “Anti-diffusive dynamics: from sub-cellular to astrophysical scales”, supported by EPSRC Grant No. EP/R014604/1, and for a Rothschild Distinguished Visiting Fellowship. KVK acknowledges support of the Department of Atomic Energy, Government of India, under Project No. RTI4001.